# The Boolean spectrum of a Grothendieck category

### v1

@@ 4. Cohomologically stable subcategories
The functor $P$ sends each differential $I^{n}\to I^{n+1}$ to zero because they are radical morphisms; see [24, Proposition 4.3.18]. Thus $H^{n}(\mathbf{R}P(X))\cong P(I^{n})$ for each $n\in\mathbb{Z}$. ∎
@@
For a Grothendieck category $\mathcal{A}$ we have alreday considered the right derived functor of the canonical functor $P\colon\mathcal{A}\to\operatorname{Spec}\mathcal{A}$. The restriction $\operatorname{Inj}\mathcal{A}\to(\operatorname{Inj}\mathcal{A})/\operatorname{Rad}(\operatorname{Inj}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Spec}\mathcal{A}$ induces an exact functor
@@
where $\mathbf{K}(\operatorname{Inj}\mathcal{A})$ denotes the category of complexes in $\operatorname{Inj}\mathcal{A}$ up to homotopy. Its computation for any $X\in\mathbf{K}(\operatorname{Inj}\mathcal{A})$ is very explicit. By removing non-zero contractible summands we may assume that the complex $X$ is homotopically minimal so that all differentials are radical morphisms; see [24, Proposition 4.3.18]. Then $P(X)$ has zero differentials.
@@ 5. Exactly definable categories
Let $\mathcal{A}$ be a Grothendieck category that is locally coherent. Thus $\mathcal{A}$ has a generating set of finitely presented objects and the full subcategory $\operatorname{fp}\mathcal{A}$ of finitely presented objects is abelian. An object $X\in\mathcal{A}$ is fp-injective if $\operatorname{Ext}^{1}(F,X)=0$ for all $F\in\operatorname{fp}\mathcal{A}$. Let $\mathcal{C}:=\operatorname{Fpinj}\mathcal{A}$ denote the full subcategory of fp-injective objects in $\mathcal{A}$. This is an exact subcategory of $\mathcal{A}$ that is closed under filtered colimits and products. Note that the exact structure of $\mathcal{C}$ is intrinsic: a sequence $0\to X\to Y\to Z\to 0$ in $\mathcal{C}$ is exact if and only if it is pure-exact, so a filtered colimit of split exact sequences. Moreover, $\mathcal{C}$ admits pure-injective envelopes, which identify with the injective envelopes in the ambient category $\mathcal{A}$.
@@
A category of the form $\operatorname{Fpinj}\mathcal{A}$ for a locally coherent category $\mathcal{A}$ is called exactly definable; it has filtered colimits, products and an intrinsic exact structure given by the pure-exact sequences.
Let $\mathcal{C}$ be an exactly definable category, given by a locally coherent category $\mathcal{A}$. Then $\mathcal{A}$ is uniqueley determined by $\mathcal{C}$, because the category of pure-injective objects in $\mathcal{C}$ identifies with the category of injectve objects in $\mathcal{A}$, and any abelian category with enough injective objects is uniquely determined by its category of injective objects. We write $\mathbf{P}(\mathcal{C})$ for $\mathcal{A}$ and call it the purity category of $\mathcal{C}$.
@@
This goes back to Gruson and Jensen [15]; see also [4, §3] or [24, 12.4].
@@
for some Serre subcategory $\mathcal{S}\subseteq\operatorname{fp}\mathcal{A}$; see [24, 12.2] for details. From the above description it follows that $\mathcal{D}$ is a thick subcategory. Moreover, $\mathcal{D}$ is closed under pure-injective envelopes; see [24, Proposition 12.2.8]. We generalise the definition of a definable subcategory in two directions, following the Definitions 3.1 and 4.3.
@@
identifies the definable subcategories of $\mathcal{C}$ with the Ziegler closed subsets of $\operatorname{Ind}\mathcal{C}$. The inverse map sends a Ziegler closed subset $\mathcal{U}$ to the definable subcategory generated by $\mathcal{U}$, which consists of all pure subobjects of products of objects in $\mathcal{U}$. This correspondence is due to Crawley-Boevey for module categories [5, 2.5] and we refer to [24, 12.2] for the general version.
@@
For instance, it is shown that any pure-injective module $X$ admits an essentially unique decomposition $X=X_{\mathrm{d}}\oplus X_{\mathrm{c}}$ into a discrete and a continuous part [17, Corollary 8.28]. This reflects the decomposition
@@
of the spectral category. For a more detailed treatment of decompositions of pure-injective modules we refer to [9]. Unfortunately, not much seems to be known about the continuous part of the spectral category, except that one knows for various classes of rings when the continuous part vanishes. On the other hand, superdecomposable modules have been constructed more or less explicitly for several rings, but these constructions do not provide much insight into the ways they decompose; see [31] for a survey. In any case, it is clear from Proposition 3.6 that the Boolean lattice $\mathbf{L}(\operatorname{PSpec}\Lambda)$ controls the decompositions of pure-injective $\Lambda$-modules.
@@
(5) $\Rightarrow$ (2): When $\Lambda$ is of finite representation type, then every $\Lambda$-module is endofinite, and in particular pure-injective [24, Theorem 13.2.10]. Let $\mathcal{D}$ be a pure-essentially closed subcategory. Then it follows that $\mathcal{D}$ is closed under pure subobjects and pure quotients. In particular, $\mathcal{D}$ is closed under filtered colimits because any filtered colimit of objects in $\mathcal{D}$ is a pure quotient of a coproduct of objects in $\mathcal{D}$. Finite representation type implies that every $\Lambda$-module decomposes into a coproduct of indecomposables, and every product of objects in $\mathcal{D}$ decomposes into indecomposables from $\mathcal{D}$, because only finitely many indecomposables are involved. This follows again from endofiniteness. Thus $\mathcal{D}$ is closed under products. ∎
@@
Beyond finite representation type the correspondences for pure-essentially closed and pure-cohomologically stable subcategories (Theorems 5.6 and 5.7) seem to be new. In particular, Proposition 5.10 shows that the extension of the Ziegler spectrum yields additional information whenever the ring is not of finite representation type. This is illustrated by looking at the category of abelian groups $\mathrm{Ab}=\operatorname{Mod}\mathbb{Z}$. Consider the Prüfer group $\mathbb{Z}(p^{\infty})$ given by a prime $p$. The direct sums of copies of $\mathbb{Z}(p^{\infty})$ form a subcategory of $\mathrm{Ab}$ which is pure-essentially closed and pure-cohomologically stable, but not definable, because it is not closed under products. More precisely, the rationals $\mathbb{Q}$ arise as direct summand of a product of copies of $\mathbb{Z}(p^{\infty})$, and then $\{\mathbb{Z}(p^{\infty}),\mathbb{Q}\}$ is the Ziegler closure of $\mathbb{Z}(p^{\infty})$.
@@ Acknowledgements
It is a pleasure to thank Greg Stevenson for several discussions and helpful comments on this work. The work was supported by the Deutsche Forschungsgemeinschaft (Project-ID 491392403 – TRR 358).

### v2

@@ 5. Cohomologically stable subcategories
The functor $P$ sends each differential $I^{n}\to I^{n+1}$ to zero because they are radical morphisms; see [27, Proposition 4.3.18]. Thus $H^{n}(\mathbf{R}P(X))\cong P(I^{n})$ for each $n\in\mathbb{Z}$. ∎
@@
For a Grothendieck category $\mathcal{A}$ we have already considered the right derived functor of the canonical functor $P\colon\mathcal{A}\to\operatorname{Spec}\mathcal{A}$. The restriction $\operatorname{Inj}\mathcal{A}\to(\operatorname{Inj}\mathcal{A})/\operatorname{Rad}(\operatorname{Inj}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Spec}\mathcal{A}$ induces an exact functor
@@
where $\mathbf{K}(\operatorname{Inj}\mathcal{A})$ denotes the category of complexes in $\operatorname{Inj}\mathcal{A}$ up to homotopy. Its computation for any $X\in\mathbf{K}(\operatorname{Inj}\mathcal{A})$ is very explicit. By removing non-zero contractible summands we may assume that the complex $X$ is homotopically minimal so that all differentials are radical morphisms; see [27, Proposition 4.3.18]. Then $P(X)$ has zero differentials.
@@ 6. Exactly definable categories
Let $\mathcal{A}$ be a Grothendieck category that is locally coherent. Thus $\mathcal{A}$ has a generating set of finitely presented objects and the full subcategory $\operatorname{fp}\mathcal{A}$ of finitely presented objects is abelian. An object $X\in\mathcal{A}$ is fp-injective if $\operatorname{Ext}^{1}(F,X)=0$ for all $F\in\operatorname{fp}\mathcal{A}$. Let $\mathcal{C}:=\operatorname{Fpinj}\mathcal{A}$ denote the full subcategory of fp-injective objects in $\mathcal{A}$. This is an exact subcategory of $\mathcal{A}$ that is closed under filtered colimits and products. In fact, the exact structure of $\mathcal{C}$ is intrinsic, because it is the smallest exact structure on $\mathcal{C}$ such that the exact sequences are closed under filtered colimits. [^4] I am grateful to Kevin Schlegel for suggesting the argument: For any object $X\in\mathcal{C}$ there is a reduced power $\bar{X}=\prod_{I}X/\mathcal{F}$ (given by an appropriate ultrafilter $\mathcal{F}$ on some sufficiently large set $I$) which is injective in $\mathcal{A}$; see [32, 4.2.5]. The canonical morphism $X\to\bar{X}$ is by construction a filtered colimit of split monos, and it factors through any mono $X\to Y$ since $\bar{X}$ is injective. Thus any mono $X\to Y$ is an admissible mono with respect to any exact structure that is closed under filtered colimits. The exact sequences in $\mathcal{C}$ are called pure-exact, and $\mathcal{C}$ admits pure-injective envelopes which identify with the injective envelopes in the ambient category $\mathcal{A}$.
@@
A category that is equivalent to one of the form $\operatorname{Fpinj}\mathcal{A}$ for a locally coherent category $\mathcal{A}$ is called exactly definable; it has filtered colimits, products and an intrinsic exact structure given by the pure-exact sequences.
Let $\mathcal{C}$ be an exactly definable category, given by a locally coherent category $\mathcal{A}$. Then $\mathcal{A}$ is uniquely determined by $\mathcal{C}$, because the category of pure-injective objects in $\mathcal{C}$ identifies with the category of injective objects in $\mathcal{A}$, and any abelian category with enough injective objects is uniquely determined by its category of injective objects. We write $\mathbf{P}(\mathcal{C})$ for $\mathcal{A}$ and call it the purity category of $\mathcal{C}$.
@@
This goes back to Gruson and Jensen [18]; see also [6, §3] or [27, 12.4].
@@
for some Serre subcategory $\mathcal{S}\subseteq\operatorname{fp}\mathcal{A}$; see [27, 12.2] for details. From the above description it follows that $\mathcal{D}$ is a thick subcategory. Moreover, $\mathcal{D}$ is closed under pure-injective envelopes; see [27, Proposition 12.2.8]. We generalise the definition of a definable subcategory in two directions, following the Definitions 3.1 and 5.3.
@@
identifies the definable subcategories of $\mathcal{C}$ with the Ziegler closed subsets of $\operatorname{Ind}\mathcal{C}$. The inverse map sends a Ziegler closed subset $\mathcal{U}$ to the definable subcategory generated by $\mathcal{U}$, which consists of all pure subobjects of products of objects in $\mathcal{U}$. This correspondence is due to Crawley-Boevey for module categories [7, 2.5] and we refer to [27, 12.2] for the general version.
@@
For instance, it is shown that any pure-injective module $X$ admits an essentially unique decomposition $X=X_{\mathrm{d}}\oplus X_{\mathrm{c}}$ into a discrete and a continuous part [20, Corollary 8.28]. This reflects the decomposition
@@
of the spectral category. For a more detailed treatment of decompositions of pure-injective modules we refer to [11]. Unfortunately, not much seems to be known about the continuous part of the spectral category, except that one knows for various classes of rings when the continuous part vanishes. On the other hand, superdecomposable modules have been constructed more or less explicitly for several rings, but these constructions do not provide much insight into the ways they decompose; see [34] for a survey. In any case, the discussion in §4 explains how the Boolean lattice $\mathbf{L}(\operatorname{PSpec}\Lambda)$ controls the decompositions of pure-injective $\Lambda$-modules; see also Example 2.12.
We characterise the absence of superdecomposable modules and may ask which Boolean lattices arise from representations of finite dimensional algebras, following ideas from [4].
@@
Problem 6.10.
Let $\Lambda$ be a finite dimensional algebra of strictly wild representation type and $L$ a complete Boolean lattice. Is there a pure-injective $\Lambda$-module $X$ such that its lattice of direct summands $\mathbf{D}(X)$ is isomorphic to $L$?
@@
(5) $\Rightarrow$ (2): When $\Lambda$ is of finite representation type, then every $\Lambda$-module is endofinite, and in particular pure-injective [27, Theorem 13.2.10]. Let $\mathcal{D}$ be a pure-essentially closed subcategory. Then it follows that $\mathcal{D}$ is closed under pure subobjects and pure quotients. In particular, $\mathcal{D}$ is closed under filtered colimits because any filtered colimit of objects in $\mathcal{D}$ is a pure quotient of a coproduct of objects in $\mathcal{D}$. Finite representation type implies that every $\Lambda$-module decomposes into a coproduct of indecomposables, and every product of objects in $\mathcal{D}$ decomposes into indecomposables from $\mathcal{D}$, because only finitely many indecomposables are involved. This follows again from endofiniteness. Thus $\mathcal{D}$ is closed under products. ∎
@@
Beyond finite representation type the correspondences for pure-essentially closed and pure-cohomologically stable subcategories (Theorems 6.6 and 6.7) seem to be new. In particular, Proposition 6.11 shows that the extension of the Ziegler spectrum yields additional information whenever the ring is not of finite representation type. This is illustrated by looking at the category of abelian groups $\mathrm{Ab}=\operatorname{Mod}\mathbb{Z}$. Consider the Prüfer group $\mathbb{Z}(p^{\infty})$ given by a prime $p$. The direct sums of copies of $\mathbb{Z}(p^{\infty})$ form a subcategory of $\mathrm{Ab}$ which is pure-essentially closed and pure-cohomologically stable, but not definable, because it is not closed under products. More precisely, the rationals $\mathbb{Q}$ arise as direct summand of a product of copies of $\mathbb{Z}(p^{\infty})$, and then $\{\mathbb{Z}(p^{\infty}),\mathbb{Q}\}$ is the Ziegler closure of $\mathbb{Z}(p^{\infty})$.
@@ Acknowledgements
It is a pleasure to thank Greg Stevenson for several discussions and helpful comments on this work. Also, I am grateful to Ken Goodearl and Jan Šťovíček for their valuable input. The work was supported by the Deutsche Forschungsgemeinschaft (Project-ID 491392403 – TRR 358).